Perimeter of rectangle is 90. One side is 6x and one end is 2x. What is the length of the longest side of the rectangle
Answer by asuar010(338)

First set up the equation for the perimeter of a rectangle:
2*(2x+6x)=90 then 2*(8x)=90 then 16x=90 then x=5.625 so the longest side is
6*5.625=33.75